OpenStudy (96mertcank96):

(5/8)-(2/3)(6-(1/2))+(3/4)=?????

OpenStudy (96mertcank96):

\[\frac{ 5 }{ 6 } -\frac{ 2 }{ 3 }\left( 6-\frac{ 1 }{ 2 } \right)+\frac{ 3 }{ 4} =?\]

OpenStudy (anonymous):

If we use PEMDAS, we begin with the parentheses. So... 6-(1/2)= (12/2) - (1/2) = (11/2) Thus, we now have... (5/6)- (2/3)*(11/2) + (3/4) Now we will multiply (2/3) and (11/2), and... (2/3) * (11/2) = (2*11)/(3*2) = (22/6), so... (5/6) - (22/6) + (3/4). Now we will finish by going from left to right. (5/6) - (22/6) = (5-22)/6 = (-17/6) (-17/6) + (3/4) = ? We need common denominators in order to add and subtract fractions, so we find a number that 4 and 6 both go into. Let's use 12, because 2*6 = 12 and 3*4=12. We must also multiply the numerators by these numbers. So we multiply -17 by 2 and 3 by 3, so... ((-17*2)/12) + ((3*3)/12) = (-34/12) + (9/12) = (-34 + 9)/12 = (-25/12)